OpenStudy (marigirl):

Find the values of t when the function defined by x=4t^3+1 and y=3t^2-7t is concave up.

OpenStudy (marigirl):

\[x=4t^3+1 \] \[\frac{ dx }{ dt }=12t^2\] and then \[y=3t^2-7t\] \[\frac{ dy }{ dt }=6t-7\] now to obtain dy/dx \[\frac{ dy }{ dx }=\frac{ 6t-7 }{ 12t^2 }\]

OpenStudy (marigirl):

now how do i go about looking for the values where it concave up? I have to take the second derivative but what is the x value we are after?

OpenStudy (anonymous):

you need to find values of t such that d^2y/dx^2 > 0

OpenStudy (marigirl):

but i cant do d^2y/dx^2 because my equation is written as t as a variable

OpenStudy (anonymous):

you can use differential notation to derive a formula

OpenStudy (marigirl):

can u plz show me

OpenStudy (marigirl):

\[d^2y/dx^2=\frac{ -72t^2+168t }{ (12t^2)^2 }\]

OpenStudy (anonymous):

$$ \large{\frac{d^2y}{dx^2 } =\frac{d}{dx}\left( \frac{dy/dt}{dx/dt} \right) = \frac{d}{dt} \left( \frac{dy/dt}{dx/dt} \right) \frac{dt}{dx} \\ = \frac{\Large \frac{d}{dt} \left( \frac{dy/dt}{dx/dt} \right) }{dx/dt} }$$ verify that all the differentials cancel and work out

OpenStudy (marigirl):

so then essentially what i need do it divide the above second differential by 12t^2 because that is dx/dt

OpenStudy (anonymous):

right. take the derivative of (6t - 7) / 12t^2 , and then divide by 12t^2

OpenStudy (marigirl):

do you think i need to go and make it greater than zero as well?

OpenStudy (anonymous):

we need the final dy^2 / dx^2 in terms of t, then find t where its greater than zero

OpenStudy (anonymous):

we could always explicitly solve for y in terms of x it might be easier x=4t^3 + 1 y = 3t^2 -7t solve for t in first equation t = ((x - 1)/4 ) ^(1/3) substitute into second equation y = 3 ( ((x - 1)/4 ) ^(1/3) ] ^2 - 7( (x - 1)/4 ) ^(1/3) y = 3( (x-1)/4 )^(2/3) - 7(( x-1)/4)^(1/3) we can see they are equivalent here https://www.desmos.com/calculator/6wh7bsuwt8

OpenStudy (anonymous):

It looks like using t is easier. I get dy^2 / dx^2 = (7 - 3t) / ( 75 t^5 ) solve this greater than zero Test the real number line , using t = 0 and t = 7/3 as your 'critical points' ------ 0 +++++ 7/3 -------

OpenStudy (anonymous):

it is concave up between t = 0 and t = 7/3 the hard part is getting the expression for the second derivative

OpenStudy (marigirl):

cool i got that same answer too

OpenStudy (marigirl):

thanks heaps

OpenStudy (anonymous):

Your welcome :)